For a particular reaction at 157.9 °C, Δ?=−346.08 kJ/mol, and Δ?=996.11 J/(mol⋅K) . Calculate ΔG for this reaction at −46.1 °C.

Δ?=
Answer #1

using AG and AS values, we find AH We know that AG= AH-TAS AH = AG + TAS = -346.08 kJ/mol + (431.05 K * 996.11 J/(mol-K)= 83.

AG = - RTINK = -8.314 J/(mol.K) * 227.05 K * In(7.42x 1032 ) = - 142873.6 J/mol = – 142.8736 kJ/mol AGºat – 46.1 °C is - 142.